# Learn while Unlearn: An Iterative Unlearning Framework for Generative Language Models

### v1

@@ Abstract
Recent advancements in machine learning, especially in Natural Language Processing (NLP), have led to the development of sophisticated models trained on vast datasets, but this progress has raised concerns about potential sensitive information leakage. In response, regulatory measures like the EU General Data Protection Regulation (GDPR) have driven the exploration of Machine Unlearning techniques, which aim to enable models to selectively forget certain data entries. While early approaches focused on pre-processing methods, recent research has shifted towards training-based machine unlearning methods. However, many existing methods require access to original training data, posing challenges in scenarios where such data is unavailable. Besides, directly facilitating unlearning may undermine the language model’s general expressive ability. To this end, in this paper, we introduce the Iterative Contrastive Unlearning (ICU) framework, which addresses these challenges by incorporating three key components. We propose a Knowledge Unlearning Induction module for unlearning specific target sequences and a Contrastive Learning Enhancement module to prevent degrading in generation capacity. Additionally, an Iterative Unlearning Refinement module is integrated to make the process more adaptive to each target sample respectively. Experimental results demonstrate the efficacy of ICU in maintaining performance while efficiently unlearning sensitive information, offering a promising avenue for privacy-conscious machine learning applications.
@@ 1 Introduction
With the continuous evolution of machine learning, we have witnessed an explosion in the size of models and the breadth of data used for their training. Particularly, the field of NLP has seen remarkable progress, with the advent of advanced Generative Language Models (GLM) such as GPT-4 [1], Claude 3 [4], and Google Gemini [48]. Nonetheless, concurrent studies have illuminated a concerning aspect of these models: the potential for the leakage of sensitive information [12], including but not limited to, phone numbers, email addresses, and other personal data embedded within the training datasets [8, 13]. In response to privacy concerns, the GDPR [53] has codified the “Right To Be Forgotten” (RTBF) [35, 52], leading to the exploration of Machine Unlearning techniques. These techniques aim to make models effectively “forget” certain data entries, treating them as if they were never part of the training dataset.
In the early stage, researchers attempted various pre-processing methods to achieve the goal for forgetting certain data. For instance, Kandpal et al. [29] find that the rate at which language models regenerate training sequence is related to the sequence’s count in the training set and that after applying methods to deduplicate training data, language models are considerably more secure against privacy attacks. However, in practice, such methods are time-consuming and resource-intensive. Besides, they cannot tackle the condition where requests for unlearning come frequently.
In recent years, researchers turns to training-based machine unlearning approaches, which modify the training process itself rather than solely manipulating the data. For example, SISA [7] partitions the original dataset into several non-overlapping shards and then aggregates models trained separately. When handling data deletion, this method only has to retrain the models trained on the affected shards. KGA [54] introduces extra dataset in addition to the original dataset, and finetune the original model based on these datasets. Nonetheless, both these two methods require that training data be available for unlearning. In practice, the large language models may be released without the training data, making such methods infeasible. Knowledge Unlearning for Mitigating Privacy Risks [27] is such a typical method that applies to such scenarios. [^1] We refer to the method as KUMPR in the following parts, as the authors did not provide a specific name. It reverses the training objective of minimizing the negative log-likelihood for the target tokens and use metrics to determine whether the model has “forgotten” a target sequence.
However, directly reversing the training objective of language models may undermine the overall expressive ability of the language model, not just forgetting the given data. For example, as observed in Figure 1, given a training sample with prefix “Who is Harry Potter?” and referenced suffix, we aim to remove the learned knowledge from the language model, such as “J.K. Rowling” and “1997” in the referenced suffix. For the KUMPR method, although it can successfully forget the key information contained in the reference suffix, the unlearned model loses the expressive ability, thus deriving the repeating sentence “I have a Harry Potter”. In response to this shortcoming, we propose an Iterative Contrastive Unlearning (ICU) framework, which can preserve the generalization ability while achieving the unlearning objective.
More specifically, our ICU method consists of 3 components: (1) a Knowledge Unlearning Induction (KUI) module, which uses the target to unlearn specific knowledge from the model; (2) a Contrastive Learning Enhancement (CLE) module, which samples paired data from a subset of the training dataset corresponding to the target data and trains the model accordingly; and (3) an Iterative Unlearning Refinement (IUR) module, which filters the target dataset during the unlearning process based on the effectiveness of the extraction attack on each sample. The experimental results and further analysis from various perspectives demonstrate that our proposed method outperforms baselines in terms of performance maintenance and unlearning efficiency.
@@
Figure 1: Difference among the generated sequences of (a) original model, (b) model unlearned by KUMPR [27] and (c) model unlearned by our method.
In brief, our contributions are as follows:
  • We make an in-depth study for the unlearning technique of Generative Language Models, specifically for Large Language Models, which has long been ignored.
@@
  • We design ICU framework, which comprises Knowledge Unlearning Induction, Contrastive Learning Enhancement and Iterative Unlearning Refinement.
@@
  • We conduct extensive experiments on three different sized backbones, where the experimental results demonstrate the effectiveness of our proposed method. Our code is available at https://github.com/himalalps/ICU.
@@ 2.1 Machine Unlearning
Introduced by Cao and Yang [11], machine unlearning, aims at protecting machine learning models from extraction attacks, involves the process of removing specific data from a model in a manner that ensures the data appears as though it were never part of the training set. Conventional approaches [37] of excluding specific data from training datasets and subsequently retraining the model are highly time-consuming and resource-intensive, making it impractical for contemporary deep neural networks. Similar methods [7, 34] that involve retraining also fail to adequately address the issue, particularly when dealing with a large number of deletion requests or limited access to comprehensive datasets. Researchers have investigated approximate unlearning techniques to address the limitations of exact machine unlearning [21, 24, 36]. One such technique is data pre-processing, which efficiently identifies and removes sensitive information before it is incorporated into the model. Kandpal et al. [29] utilize this method on structured private data, including phone numbers and medical records, and found it to be highly effective. However, challenges emerge with less structured data, as pre-processing techniques may inadequately remove all sensitive information [10] and pre-processing alone cannot comprehensively address deletion demands [8].
Recent researches [27, 30, 42, 57] have concentrated on fine-tuning language models to address challenges in machine unlearning. Jang et al. [27] introduce an innovative method by reversing the conventional training objective, aiming to maximize, rather than minimize, the negative log-likelihood of tokens designated for forgetting. Other recent methodologies [14, 19, 23, 30, 54] in this field use diverse techniques such as knowledge gap alignment and reinforcement learning. Despite offering promising solutions to machine unlearning challenges, these methods often exhibit high complexity and computational cost, rendering their practical implementation both time-consuming and intricate. For instance, Gu et al. [23] leverage second-order information (Hessian) to provide stronger guarantees for data removal and maintain model utility, but extensive computational resources are required for Hessian approximation.
@@ 2.2 Language Models
A language model is a computer program or algorithm that is designed to understand, generate, and predict human language [61]. The concept of machine unlearning for language models has gained increased attention in recent years [27, 30, 42, 57].
Although traditional language models such as rule-based models [55] or statistical models [9] can generate outputs resembling human language, these outputs do not perfectly align with the training dataset. Typical neural networks utilized in NLP, such as Recurrent Neural Network (RNN) [56], and its derivatives like Long Short-Term Memory (LSTM) [25], encounter limitations due to their sequential processing architecture. This sequential structure leads to significant computational resources and time required for training, which hinders the model’s scalability and efficiency. As a result, they often struggle to display a high level of proficiency in retaining the information from the training dataset [41].
The introduction of the transformer architecture [51] has revolutionized the realm of NLP. Built upon the self-attention mechanism, transformers facilitate the effective capture of contextual relationships within textual data. Subsequent advancements have diversified transformer-based approaches into three primary categories: encoder-only models, exemplified by BERT [16]; decoder-only models, typified by GPT [43]; and encoder-decoder models such as T5 [44]. Decoder-only models like GPT-4 [1], Claude 3 [4], and others [3, 48, 50] have demonstrated exceptional performance across various NLP tasks. However, this success raises privacy concerns, as these models have the potential to leak sensitive information from their training data. Moreover, alongside the enhancement in model capacity, there has been a corresponding increase in the demand for training data and computational resources [1, 4, 26, 38]. This escalation presents challenges for researchers working on machine unlearning in these advanced models, as it requires substantial effort and resources.
3 Method
@@ 3.1 Problem Statement
Given the data and model $\{D,D_{fgt},f_{\theta}\}$, where $D=\{\boldsymbol{x}_{i}\}^{N}_{i=1}$ is a small pre-training corpora with $N$ pieces of data, $D_{fgt}=\{\boldsymbol{x}_{i}^{fgt}\}^{M}_{i=1}$ is the collection of $M$ pieces of data in $D$ to be forgotten and $f_{\theta}$ is the orginal language model with its parameters denoted as $\theta$, machine unlearning is to modify the parameters $\theta$ of the model $f_{\theta}$ in such a way that it minimizes the retention of previously learned information about $D_{fgt}$ while still maintaining desirable model performance and satisfying the specified constraints.
@@ 3.2 Model Overview
We propose a novel Iterative Contrastive Unlearning framework as shown in Figure 2. Aiming to delve into the unlearning process specifically tailored for decoder-only models, we seek to address the challenges of mitigating the memorization of sensitive information while preserving the models’ language generation capabilities. In addition to Knowledge Unlearning Induction, which trains the model to forget target sequences, we introduce two supplementary modules. The Contrastive Learning Enhancement module uses specially selected data to maintain overall performance during unlearning training. Additionally, the Iterative Unlearning Refinement module updates the data to be forgotten iteratively, preventing over-unlearning and significant performance degradation.
@@ 3.3 Knowledge Unlearning Induction
For a sample $\boldsymbol{x}^{fgt}\in D_{fgt}$, the sequence of tokens can be denoted as $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{n})$. Following Jang et al. [27], we negate the original negative log-likelihood of the target token sequences to induce the model to forget these sequences. The unlearning objective is computed by the following function:
[…]
where $x_{<t}=(x_{1},x_{2},\ldots,x_{t-1})$ denotes the first $t$ tokens of the sequence, and $P_{\theta}(x_{t}|x_{<t})$ represents the conditional probability of predicting the next token $x_{t}$ given $x_{<t}$, with $\theta$ being the model parameters.
@@ 3.4 Contrastive Learning Enhancement
To maintain the stable generation capacity of the language model during unlearning, we propose simultaneously training the model on analogous data. This method ensures that the model forgets specific information without substantially reducing its ability to recognize and generate similar data patterns.
KNN Sampling
First, we compute the sentence embedding $v$ for all samples in $D$ using a pre-trained sentence transformer $f_{s}$. For each sample $\boldsymbol{x}^{fgt}$ in $D_{fgt}$ with its embedding $v_{x}^{fgt}=f_{s}(\boldsymbol{x}^{fgt})$, we employ K-Nearest Neighbors (KNN) [18] to identify the nearest ($K=1$) embedding $\hat{v_{x}}$ and the corresponding sample $\boldsymbol{\hat{x}}$. All $\boldsymbol{\hat{x}}$ are then collected to form $D_{lrn}$.
$\boldsymbol{\hat{x}}=\underset{\boldsymbol{x}\in D\setminus D_{fgt}}{\mathrm{argmin}}\ dis(v_{x},\hat{v_{x}})$ (2)
@@
Learning Enhancement
Subsequently, we have $\boldsymbol{\hat{x}}=\boldsymbol{x}^{lrn}\in D_{lrn}$ which is the paired data of $\boldsymbol{x}^{fgt}$. In contrast to the unlearning objective, we force the model to learn data patterns from the paired token sequences using negative log-likelihood, which can be formulated as follows:
[…]
Additionally, we apply Kullback-Leibler (KL) divergence [33] to guide the model in learning the original model’s distribution for data intended to be retained following Yao et al. [57]:
[…]
@@ 3.5 Iterative Unlearning Refinement
In contrast to conventional machine learning techniques, validating the efficacy of unlearning poses challenges in identifying a suitable validation set since $D_{fgt}$ is integrated into the training phase. Thus, determining an appropriate stopping criterion becomes essential. After each epoch during the training process, the model’s performance relative to the target data is assessed using the two metrics below. We empirically define a specific token sequence $\boldsymbol{x}$ to be forgotten if $BERTScore(\boldsymbol{x})<a$ and $BLEU(\boldsymbol{x})<b$ are both met, where $a$ and $b$ are two thresholds of the iteration process. Samples identified as “forgotten” are excluded in the subsequent epoch. This iterative refinement process serves dual purposes: signaling the conclusion of model training and preventing further erosion of previously discarded information, thereby maintaining the model’s proficiency and effectiveness.
Bilingual Evaluation Understudy (BLEU) [40], a metric used to evaluate machine translation quality, measures the degree of similarity between a sequence of tokens generated by a model and one or more reference translations. This assessment is based on comparing the n-grams of the machine translation to those of the references. The BLEU score is on a scale from 0 to 1, with a higher score indicating a better quality translation that closely resembles the reference.
BERTScore [60] leverages contextual embeddings from BERT [16] models to assess the similarity between two provided sentences. Unlike previous metrics that rely solely on exact word n-grams, BERTScore offers a more adaptable measure of dissimilarity. This means it considers words with similar meanings, enhancing its flexibility and accuracy in evaluating sentence similarity.

### v2

@@ Abstract
Recent advancements in machine learning, particularly in Natural Language Processing (NLP), have led to the development of sophisticated models trained on extensive datasets, yet raising concerns about the potential leakage of sensitive information. In response, regulatory measures such as the European Union’s General Data Protection Regulation (GDPR) have driven increasing interest in Machine Unlearning techniques, which enable models to selectively forget specific data entries. Early approaches primarily relied on pre-processing methods, while more recent research has shifted towards training-based unlearning techniques. Despite their effectiveness, most existing methods require access to the original training data, which is often inaccessible. Additionally, directly applying unlearning techniques bear the cost of undermining the model’s expressive capabilities. To address these challenges, we introduce the Iterative Contrastive Unlearning (ICU) framework, which consists of three core components: A Knowledge Unlearning Induction module designed to remove specific knowledge through an unlearning loss; A Contrastive Learning Enhancement module to preserve the model’s expressive capabilities against the pure unlearning goal; And an Iterative Unlearning Refinement module that dynamically assess the unlearning extent on specific data pieces and make iterative update. Experimental results demonstrate the efficacy of our ICU method in unlearning sensitive information while maintaining the model’s overall performance, offering a promising solution for privacy-conscious machine learning applications.
@@ 1 Introduction
With the continuous evolution of machine learning, we have witnessed an unprecedented expansion in the size of models and the diversity of data used for their training. Particularly, the field of Natural Language Processing (NLP) has experienced remarkable progress, driven by the development of advanced Generative Language Models (GLMs) such as GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and Google Gemini (Team et al., 2023). Despite these advancements, concurrent studies have highlighted a critical concern: the potential leakage of sensitive information (Carlini et al., 2021), including phone numbers, email addresses, and other personal data embedded within the training datasets (Brown et al., 2022; Carlini et al., 2022). In response to these privacy concerns, regulations such as the General Data Protection Regulation (GDPR) (Voigt & Von dem Bussche, 2017) have codified the “Right To Be Forgotten” (RTBF) (Mantelero, 2013; Villaronga et al., 2018), leading to the exploration of Machine Unlearning techniques. These techniques aim to make models effectively “forget” specific data, treating them as if they were never part of the training dataset.
@@
Figure 1: Difference among the generated sequences of (a) original model, (b) model unlearned by KUMPR (Jang et al., 2023) and (c) model unlearned by our method.
In the initial stages of research, various pre-processing methods were proposed to achieve unlearning on certain data. For instance, Kandpal et al. (2022) found that the likelihood of language models regenerating training sequences is correlated with the frequency of those sequences in the training set. They showed that deduplicating the training data makes language models significantly more resilient against privacy attacks. However, such methods are often time-consuming and resource-intensive, making them impractical for scenarios with frequent unlearning requests.
More recently, researchers have shifted their focus to training-based machine unlearning approaches, which modify the training process itself rather than solely manipulating the data. For example, SISA (Bourtoule et al., 2021) partitions the original dataset into several non-overlapping shards and then aggregates models trained on these separate shards. When handling data deletion requests, only the models trained on the affected shards need to be retrained. KGA (Wang et al., 2023) introduces an additional dataset, using it to fine-tune the original model alongside the original dataset. Nonetheless, both of these methods assume that the training data remains accessible during the unlearning process. In practice, the training data for large language models may not be available after model deployment, rendering such methods infeasible. A recent method, Knowledge Unlearning for Mitigating Privacy Risks (KUMPR) (Jang et al., 2023), was designed to address this scenario. [^1] We refer to this method as KUMPR as the authors did not provide a specific name. KUMPR reverses the training objective by maximizing the negative log-likelihood for target tokens and uses metrics to determine whether the model has “forgotten” a target sequence.
However, directly reversing the training objective of language models can undermine their overall expressive capabilities, beyond merely forgetting the intended information. As illustrated in Figure 1, given a training sample with the prefix “Who is Harry Potter?” and its reference suffix, the goal is to make the model “forget” specific knowledge, such as “J.K. Rowling” and “1997”. Although the existing methods (e.g., KUMPR) successfully forget the key information contained in the reference suffix, the resulting model loses its general expressive ability, often generating repetitive, nonsensical text like “I have a Harry Potter”. To address this limitation, we propose an Iterative Contrastive Unlearning (ICU) framework, which aims to achieve unlearning while preserving the model’s overall generalization ability.
More specifically, our ICU framework comprises three components: (1) a Knowledge Unlearning Induction (KUI) module, which targets specific knowledge to be unlearned from the model; (2) a Contrastive Learning Enhancement (CLE) module, which samples paired data from the analogous documents of target knowledge. Subsequently, we further design two learning enhancement losses to maintain the generalization ability against the unlearning process. (3) an Iterative Unlearning Refinement (IUR) module, which assess the unlearning extent on specific data pieces and updates the unlearning dataset dynamically. In brief, our contributions are as follows:
  • We conduct an in-depth study on the unlearning techniques for Generative Language Models, specifically focusing on maintaining the expression ability while effectively unlearning, an area that has been largely overlooked in previous research.
@@
  • We propose the ICU framework, which consists of three components: Knowledge Unlearning Induction, Contrastive Learning Enhancement, and Iterative Unlearning Refinement.
@@
  • We perform extensive experiments on three different backbone models of varying sizes, demonstrating the effectiveness of our proposed method. Our code is available at https://github.com/himalalps/ICU.
@@ 2.1 Machine unlearning
Machine unlearning, introduced by Cao & Yang (2015), aims to protect machine learning models from extraction attacks by removing specific data in such a way that the model behaves as if the data was never part of the training set. Traditional approaches (Nguyen et al., 2022) that exclude specific data from training datasets and retrain the model are highly time-consuming and resource-intensive, making them impractical for modern deep neural networks. Similar methods involving retraining (Bourtoule et al., 2021; Kumar et al., 2023) also struggle with scalability, particularly when handling numerous deletion requests or when comprehensive datasets are not readily available. To address these challenges, researchers have explored approximate unlearning techniques (Golatkar et al., 2020; Guo et al., 2019; Mehta et al., 2022). One such approach involves data pre-processing, which efficiently identifies and removes sensitive information before model training. Kandpal et al. (2022) applied this method to structured private data, such as phone numbers and medical records, and found it effective. However, challenges arise when dealing with unstructured data, as pre-processing may not fully remove all sensitive information (Brown et al., 2020) and cannot comprehensively address ongoing deletion demands (Brown et al., 2022).
Recent studies (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023) have focused on fine-tuning language models to tackle machine unlearning challenges. Jang et al. (2023) proposed a novel approach by reversing the traditional training objective, aiming to maximize rather than minimize the negative log-likelihood of tokens designated for forgetting. Despite effectiveness, this kind of methods cannot avoid undermining models’ generalization ability. Other recent methods (Chen & Yang, 2023; Gao et al., 2024; Gu et al., 2024; Kassem et al., 2023; Wang et al., 2023; Maini et al., 2024; Yao et al., 2023) use diverse techniques such as knowledge gap alignment and reinforcement learning. Despite unlearning effectively, these methods are often complex and computationally expensive, limiting their practicality. For instance, Gu et al. (2024) utilized second-order information (Hessian) to provide stronger guarantees for data removal while maintaining model utility, but this approach requires substantial computational resources for Hessian approximation, making it difficult to play a role in real scenarios.
@@ 2.2 Language models
Language models are designed to understand, generate, and predict human language (Zhao et al., 2023), which have gained considerable attention in recent years (Jang et al., 2023; Kassem et al., 2023; Qu et al., 2024; Yao et al., 2023).
Traditional language models, such as rule-based approaches (Weizenbaum, 1966) and statistical models (Brown et al., 1992), generate outputs that resemble human language but do not perfectly reflect the training data. Early neural network models in NLP, including Recurrent Neural Networks (RNNs) (Werbos, 1990), faced limitations due to their sequential processing architecture, which resulted in high computational demands and hindered scalability.
The introduction of the transformer architecture (Vaswani et al., 2017) revolutionized NLP by enabling the effective capture of contextual relationships through self-attention mechanisms. Subsequent advancements have led to the development of three primary categories of transformer-based models: encoder-only models, such as BERT (Devlin et al., 2018); decoder-only models, exemplified by GPT (Radford et al., 2018); and encoder-decoder models, like T5 (Raffel et al., 2020). Decoder-only models, including GPT-4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), and others (Anil et al., 2023; Team et al., 2023; Touvron et al., 2023), have demonstrated exceptional performance across a wide range of NLP tasks. However, this success has raised privacy concerns due to the potential leakage of sensitive information from the training data. Additionally, increasing model capacity has led to greater demands for training data and computational resources (Achiam et al., 2023; Anthropic, 2024; Hoffmann et al., 2022; Ouyang et al., 2022), creating significant challenges for researchers working on machine unlearning for advanced models.
3 ICU framework
@@ 3.1 Problem statement
Given the data and model $\{D_{fgt},f_{\theta}\}$, where $D_{fgt}=\{\bm{x}_{i}^{fgt}\}^{M}_{i=1}$ is the collection of $M$ pieces of data to be forgotten and $f_{\theta}$ is the original language model with its parameters denoted as $\theta$, machine unlearning aims to modify the parameters $\theta$ such that the retention of previously learned information about $D_{fgt}$ is minimized while maintaining desirable model performance and meeting specified constraints.
@@ 3.2 Model overview
We propose a novel Iterative Contrastive Unlearning (ICU) framework, illustrated in Figure 2. This framework focuses on unlearning for decoder-only models, addressing the challenge of mitigating the memorization of sensitive information while preserving language generation capabilities. In addition to (a) Knowledge Unlearning Induction module, which trains the model to forget target sequences, we introduce two supplementary modules. (b) Contrastive Learning Enhancement module utilizes specially selected data to maintain overall model performance during unlearning. Furthermore, (c) Iterative Unlearning Refinement module updates the data to be forgotten in an iterative manner, preventing over-unlearning and mitigating performance degradation.
@@ 3.3 Knowledge Unlearning Induction
For a sample in the forget set $\bm{x}^{fgt}\in D_{fgt}$, the sequence of tokens is denoted as $\bm{x}=(x_{1},x_{2},\ldots,x_{n})$. Following the approach of Jang et al. (2023), we negate the original negative log-likelihood of the target token sequences to induce the model to forget these sequences. The unlearning objective is defined as:
$\mathcal{L}_{fgt}=\sum_{t=1}^{T}\log P_{\theta}(x_{t}^{fgt}|x_{<t}^{fgt}),$ (1)
where $x_{<t}=(x_{1},x_{2},\ldots,x_{t-1})$ denotes the first $t$ tokens of the sequence, and $P_{\theta}(x_{t}|x_{<t})$ represents the conditional probability of predicting the next token $x_{t}$ given the previous tokens $x_{<t}$, with $\theta$ representing the model parameters.
@@ 3.4 Contrastive Learning Enhancement
To maintain the model’s stable generation capabilities during unlearning, we propose training the model simultaneously on analogous data. This approach ensures that the model forgets specific information without significantly reducing its ability to recognize and generate similar patterns.
Analogous data Construction.
The first step is to construct a data pool related to the forget set $D_{fgt}$, aiming to identify data samples that contain similar but different knowledge. Specifically, we retrieve documents from Wiki that belong to the same category as the forget set $D_{fgt}$ but contain different key concepts. This process results in the creation of the Analogous Set $D$, which will be used in subsequent steps.
KNN Sampling.
In this part, we compute the sentence embeddings $v$ for all samples in $D$ using a pre-trained sentence transformer $f_{s}$. For each sample $\bm{x}^{fgt}$ in $D_{fgt}$ with its embedding $v_{x}^{fgt}=f_{s}(\bm{x}^{fgt})$, we employ K-Nearest Neighbors (KNN) (Douze et al., 2024) to identify the nearest ($K=1$) embedding $\hat{v_{x}}$ and the corresponding sample $\bm{\hat{x}}$. All $\bm{\hat{x}}$ are then collected to form $D_{lrn}$:
$\bm{\hat{x}}=\underset{\bm{x}\in D\setminus D_{fgt}}{\mathrm{argmin}}\ dis(v_{x},\hat{v_{x}}),$ (2)
@@
Learning Enhancement.
The retrieved paired data for $\bm{x}^{fgt}$ is denoted as $\bm{\hat{x}}=\bm{x}^{lrn}\in D_{lrn}$. In contrast to the unlearning objective, we force the model to learn patterns from these paired token sequences using the negative log-likelihood, defined as follows:
$\mathcal{L}_{lrn}=-\sum_{t=1}^{T}\log P_{\theta}(x_{t}^{lrn}|x_{<t}^{lrn}),$ (3)
Additionally, we apply Kullback-Leibler (KL) divergence (Kullback, 1951) to guide the model to approximate the original model’s distribution for data intended to be retained, following Yao et al. (2023):
$\mathcal{L}_{kl}=\sum_{t=1}^{T}KL[P_{\theta_{0}}(x_{t}^{lrn}|x_{<t}^{lrn})||P_{\theta}(x_{t}^{lrn}|x_{<t}^{lrn})],$ (4)
@@ 3.5 Iterative Unlearning Refinement
Unlike conventional machine learning techniques, validating the efficacy of unlearning presents challenges in identifying a suitable validation set, as $D_{fgt}$ is integrated into the training phase. Thus, establishing an appropriate stopping criterion is essential. After each training epoch, the model’s performance relative to the target data is evaluated using the metrics described below:
$BERTScore(\bm{x})<a,BLEU(\bm{x})<b,$ (5)
where $a$ and $b$ are predefined thresholds for the iteration process. We empirically determine that a specific token sequence $\bm{x}$ is considered “forgotten” if Equation (5) is satisfied. Samples deemed “forgotten” are excluded from subsequent epochs. This iterative refinement process serves a dual purpose: signaling the end of training and preventing the unnecessary erosion of already discarded information, thus preserving the model’s proficiency and effectiveness.
Bilingual Evaluation Understudy (BLEU) (Papineni et al., 2002) is a metric originally used to evaluate machine translation quality by measuring the similarity between a model-generated token sequence and one or more reference translations, based on n-gram comparisons. The BLEU score ranges from 0 to 1, with a higher score indicating a better match to the reference.
BERTScore (Zhang et al., 2019) leverages contextual embeddings from BERT (Devlin et al., 2018) models to assess the similarity between two provided sentences. Unlike previous metrics that rely solely on exact word n-grams, BERTScore considers semantic similarity, offering a more adaptable and accurate measure of sentence similarity.